# Bayesian Signal Matching for Transfer Learning in ERP-Based Brain Computer Interface

### v1

@@ Abstract
An Event-Related Potential (ERP)-based Brain-Computer Interface (BCI) Speller System assists people with disabilities communicate by decoding electroencephalogram (EEG) signals. A P300-ERP embedded in EEG signals arises in response to a rare, but relevant event (target) among a series of irrelevant events (non-target). Different machine learning methods have constructed binary classifiers to detect target events, known as calibration. Existing calibration strategy only uses data from participants themselves with lengthy training time, causing biased P300 estimation and decreasing prediction accuracy. To resolve this issue, we propose a Bayesian signal matching (BSM) framework for calibrating the EEG signals from a new participant using data from source participants. BSM specifies the joint distribution of stimulus-specific EEG signals among source participants via a Bayesian hierarchical mixture model. We apply the inference strategy: if source and new participants are similar, they share the same set of model parameters, otherwise, they keep their own sets of model parameters; we predict on the testing data using parameters of the baseline cluster directly. Our hierarchical framework can be generalized to other base classifiers with clear likelihood specifications. We demonstrate the advantages of BSM using simulations and focus on the real data analysis among participants with neuro-degenerative diseases.
Keywords: Bayesian Method, Mixture Model, Data Integration, P300, Brain-Computer Interface, Calibration-Less Framework.
@@ 1.1 Background
A Brain-Computer Interface (BCI) is a device that interprets brain activity to assist people with severe neuro-muscular diseases with normal communications. An electroencephalogram (EEG)-based BCI speller system is one of the most popular BCI applications, which enables a person to “type” words without using a physical keyboard by decoding the recorded EEG signals. The EEG brain activity has the advantages of non-invasiveness, low cost, and high temporal resolution (Niedermeyer and da Silva,, 2005).
The P300 event-related potential (ERP) BCI design, known as the P300 ERP-BCI design (Donchin et al.,, 2000), is one of the most widely used BCI frameworks and the focus of this manuscript. However, other types of ERPs exist that help interpret and classify brain activity. An ERP is a signal pattern in the EEG signals in response to an external event. The P300 ERP is a particular ERP that occurs in response to a rare, but relevant event (i.e., highlighting a group of characters on the screen) among a group of frequent, but irrelevant ones. The relevant (target) P300 ERP usually has a positive deflection in voltage with the latency (the time intervals between the onset of the event and the first response peak) around 300 ms.
@@
Under the P300 ERP-BCI design, participants wore an EEG cap, sat next to a virtual screen, and focused on a specific character to spell, denoted as the “target character.” The study applied the row-and-column paradigm (RCP) developed by (Farwell and Donchin,, 1988), which is a 6 $\times$ 6 grid of characters. Each event is either a row stimulus or a column stimulus. The order of rows and columns is random, but loops through all rows and columns every consecutive 12 stimuli, known as a sequence. Participants were asked to mentally count whenever they saw a row or column stimulus containing the target character or to do nothing. Thus, within each sequence, there were always exactly two stimuli that were supposed to elicit target P300 ERPs.
@@
The fundamental operation of the P300 ERP-BCI design is that it presents a sequence of stimuli on a virtual keyboard and that it analyzes the EEG signals in a fixed time response window post event to make a binary decision whether the participant’s brain elicits a target P300 ERP response. After each method produces binary classifier scores for each stimulus, those scores are converted into character-level probabilities. Finally, the character with the highest probability is shown on the virtual screen. Many machine learning (ML) methods have successfully constructed such binary classifiers for P300 ERPs, including stepwise linear discriminant analysis (swLDA) (Donchin et al.,, 2000), (Krusienski et al.,, 2008), support vector machine (SVM) (Kaper et al.,, 2004), independent component analysis (ICA) (Xu et al.,, 2004), linear discriminant analysis (LDA) with xDAWN filter (Rivet et al.,, 2009), convolutional neural network (CNN) (Cecotti and Graser,, 2010), and Riemannian geometry (Barachant et al.,, 2011). These methods usually concatenate multi-channel EEG signals without considering spatial dependency. Here, an EEG channel refers to an electrode placed on the scalp to achieve stable prediction accuracy. Finally, Ma et al., (2022) attempted to improve prediction accuracy and identify spatial-temporal discrepancies between target and non-target P300 responses from the generative perspective. The method also explicitly addressed the overlapping ERP issue directly.
1.2 Challenge and Existing Work
Before participants performed the free-typing sessions, they usually copied a multi-character phrase to build a participant-specific training profile. We referred to the previous two procedures as testing and calibration, respectively. Since the signal-to-noise ratio (SNR) of EEG signals was generally low, current calibration strategy simply collected data from participants themselves with a fixed but large amount of sequence replications. However, a lengthy calibration procedure might cause attention shift and mental fatigue. Participants tended to feel bored and might elicit biased target P300 ERPs, which made the calibration process inaccurate and inefficient. Therefore, the challenge became how to reduce participants’ calibration time with decent prediction accuracy for the free-typing sessions. Existing work has tackled this problem by applying the idea of transfer learning, which was originally introduced by Bozinovski and Fulgosi, in 1976, that information was extracted and stored from existing problems to solve a new but similar problem. In statistical fields, we denoted this concept as data integration (Lenzerini,, 2002). To avoid confusion, we do not distinguish among “transfer learning”, “data integration”, or “borrowing” in our study.
@@ 1.2 Challenge and Existing Work
This subsection briefly introduces existing works that have explored the transfer learning under different domains for information leveraging with applications to P300 ERP-BCIs. A recent review (Wu et al.,, 2020) on transfer learning summarized methods between 2016 and 2020 with six paradigms and applications. For P300-ERP BCI paradigm, they considered cross-session/participant and cross-device transfer learning scenarios and produced a nice table for alignment-related approaches including offline/online status, known/unknown labels, alignment framework, referencing objects, base classifiers, and computational costs. We will continue introducing the work that were not covered previously.
First, ensembles was an intuitive idea to incorporate data from other domains that combined the results of different classifiers within the same training set. Each classifier made predictions on a test set, and the results were combined with a voting process. Rakotomamonjy and Guigue, in 2008 and Johnson and Krusienski, in 2009 were the first to apply the ensemble method to P300 ERP-BCIs by averaging the outputs of multiple SVMs and swLDAs, respectively, where base binary classifier was trained on a small part of the available data. Völker et al., in 2018 and Onishi, in 2020 applied the ensemble method by averaging the outputs of multiple CNNs to visual and auditory P300 ERP-BCI datasets, respectively. Onishi and Natsume, also mentioned that ensemble methods with overlapping partitioning criterion yielded better prediction performance than the ensemble methods with a naive partitioning criterion.
@@
As an alternative solution, Xu et al., in 2015 proposed the Ensemble Learning Generic Information (ELGI), which combined the data of the new participant with the data of source participants to form a hybrid ensemble. They split the data of each source participant into target and non-target subsets. They applied the swLDA method to construct the base classifiers by combining different subsets as follows: the target and non-target subsets from the new participant, the target subset from the new participant and the non-target subset from each source participant, and the target subset from each source participant and the non-target subset from the new participant. Thus, the resulting ensemble had $(2N+1)$ base classifiers, where $N$ is the number of source participants. They further introduced the Weighted Ensemble Learning Generic Information (WELGI) (Xu et al.,, 2016) by adding weights to each base classifier. Similarly, An et al., in 2020 proposed a weighted participant-semi-independent classification method (WSSICM) for P300 ERP-based BCIs, where they used the SVM method as the base classifier. The base classifier was fit by combining the entire data of each source participant and a small portion of data of the new participant. An ad-hoc approach was applied to determine the weighted coefficients of base classifiers for participant selection. Likewise, Adair et al., in 2017 proposed the Evolved Ensemble Learning Generic Information (eELGI). The authors argued that grouping training sets by participants was not an optimal selection criterion. Instead, they developed an evolutionary algorithm by permuting datasets among source participants to form the base classifiers, which were constructed using the swLDA method.
Finally, Riemannian geometry has gained increasing attention recently due to its fast speed to converge and a natural framework to leverage information from source participants. The Riemannian geometry classifier was originally proposed by Barachant et al., in 2011. The input data for Riemannian geometry were the sample covariance matrices. The distance-based algorithm based on the Riemannian geometry was called Minimum Distance to Mean (MDM). Instead of computing the Euclidean distance between EEG signals within the Euclidean space, the MDM method computed the Riemmanian distance between sample covariance matrices on the Riemannian manifold, and predicted the class label of which “mean” covariance matrix was the closest to the new covariance matrix with respect to the Riemmannian distance. Additional studies adapted the MDM classifier to the P300 ERP-BCI design by augmenting the covariance matrix with the label-specific reference signal patterns (Congedo et al.,, 2013), (Barachant and Congedo,, 2014). This modified classifier was a combination of first- and second-order statistics and compensated for the loss of the temporal structure information with the referencing signal components.
@@
Recent studies have shown a growing interest in building the transfer learning with the Riemannian geometry. For example, Rodrigues et al., in 2018 presented a transfer learning approach to tackle the heterogeneity of EEG signals across different sessions or participants using the Riemannian procrustes analysis (RPA). Before the authors applied the MDM classifier, they applied certain affine transformations to raw participant-level covariance matrices such that the resulting covariance matrices were less heterogeneous across sessions or participants while their Riemannian distances were preserved. Li et al., in 2020 also standardized the covariance matrices across participants by applying the affine transformation with the participant-specific Riemannian geometric mean covariance matrix. Finally, Khazem et al., in 2021 proposed another transfer learning approach, denoted as Minimum Distance to Weighted Mean (MDWM). They combined estimated mean covariance matrices from source participants and the new participant by the definition of Riemannian distance. They controlled the trade-off between new and source contribution by the power parameter, but they treated them as a hyper-parameter and did not estimate it during the calibration session. Although these three methods shared the idea that they applied different transformations to the borrowed data so that the covariance matrices were robust to aggregate information from source participants, Khazem et al., emphasized that further improvement in prediction accuracy could benefit from a proper selection of source participants.
@@ 1.3 Our Contributions
To reduce the calibration time while maintaining similar prediction accuracy, we propose a Bayesian Signal Matching (BSM) method to build a participant-dependent, calibration-less framework. BSM reduces the calibration time of a new participant by borrowing data from pre-existing source participants’ pool. Instead of transforming and homogenizing source calibration data and selecting on the level of stimuli, we simply borrow data on the level of participants, which is an intuitive clustering criterion. BSM specifies the joint distribution of stimulus-specific EEG signals from all participants via a Bayesian hierarchical mixture model. Our method has advantages from both inference and prediction perspectives. First, unlike the conventional clustering approach, we specify the baseline cluster as the one that matches the new participant. The selection indicators indicates how close source calibration data resemble the new data on the level of participants. If the selection indicator is 1, we merge source participant’s data with the new participant’s data, otherwise, we keep the source participant’s data as a separate cluster. Second, we use the baseline cluster to predict on the testing data directly without refitting the model with the augmented data. Finally, our proposed hierarchical framework is flexible to generalize to other classifiers with specific likelihood specifications.
Figure 1: An illustration of proposed data integration framework. Given rich and known data from the source participants $1,2,3,\cdots$ and a small amount of training data from the new participant 0, we construct the BSM binary classifier, apply it to the free-typing datasets and convert binary classification scores to character-level probabilities with participant-specific selection indicators that demonstrates the similarity of signals between source participants and participant 00.
The paper is organized as follows: Section 2 presents the framework for the Bayesian hierarchical mixture model using the generative parametric form as the base classifier. Section 3 introduces the details for posterior inference. Sections 4 and 5 present the numerical results for simulation and real data analysis, respectively. Section 6 concludes the paper with a discussion.
@@ 2.1 Basic Notations and Problem Setup
2.2 A Bayesian Signal Matching Clustering Framework
Given sufficient training data from the source participants and a small amount of training data from the new participant, our primary goal is to reduce calibration time by integrating a proper subset of training data on the level of source participants into the augmented training data for the new participant. We made three assumptions to simplify the model representation as follows: First, we let all source participants spell the same character $\omega$ with the same sequence replication size $I$, so we drop the target character index $l$ and the source participant index $n$ from $I_{n}$. Second, we assumed that target and non-target ERP functions shared the same covariance matrix within the same cluster. Finally, we performed selection with respect to the target ERP data and we excluded non-target ERP data from the source participants’ pool. Figure 1 provides a detailed illustration of the data integration framework of proposed Bayesian signal matching method.
@@ 2.2 A Bayesian Signal Matching Clustering Framework
where $\bm{B}_{0,1},\bm{B}_{0,0}$, and $\bm{\epsilon}_{0,i,j}$ are target ERP matrix, non-target ERP matrix, and random noise matrix of the new participant, of the same size $T_{0}\times E$, respectively. For source participant $n,n>0$, we introduce a participant-specific binary indicator $Z_{n}\in\{0,1\}$ and assume that
$\operatorname{Pr}(Z_{n}=1)=\pi_{n},\$ (2)
where $\pi_{n}$ is the probability whether source participant $n$ matches the new participant. With $\{Z_{n}\}$, we consider a signal matching model for the EEG signal matrix for source participant $n$ as follows:
$\quad\quad\quad\quad\quad\bm{X}_{n,i,j}=\{\bm{B}_{0,1}Y_{n,i,j}+\bm{B}_{0,0}(1-Y_{n,i,j})\}Z_{n}\\ +\{\bm{B}_{n,1}Y_{n,i,j}+\bm{B}_{n,0}(1-Y_{n,i,j})\}(1-Z_{n})+\bm{\epsilon}_{n,i,j},\quad\quad\quad$ (3)
where $\bm{B}_{n,1},\bm{B}_{n,0}$, and $\bm{\epsilon}_{n,i,j}$ are target ERP matrix, non-target ERP matrix, and stimulus-specific random error matrix of source participant $n$, of the same size $T_{0}\times E$, respectively. Of note, (2) and (3) assume that each source participant has either has the new participant’s parameter set or their own parameter set. However, we do not borrow non-target EEG data of source participants for making statistical inferences. All $\bm{X}_{n,i,j},\bm{B}_{n,1}$, $\bm{B}_{0,0}$, and $\bm{\epsilon}_{n,i,j}$ are matrices with the same size of $E$ by $T_{0}$.
@@
$\bm{\xi}_{n,i,j}$ $=(\bm{\xi}_{n,i,j,1},\cdots,\bm{\xi}_{n,i,j,T_{0}})^{\top},$
$\bm{\varepsilon}_{n,i,j}$ $=(\bm{\varepsilon}_{n,i,j,1},\cdots,\bm{\varepsilon}_{n,i,j,E}),$
where $\bm{\xi}_{n,i,j,t},(t=1,\cdots,T_{0})$ is a vector of spatial random effects for $E$ channels at time point $t$, and $\bm{\varepsilon}_{n,i,j,e},(e=1,\cdots,E)$ is a vector of temporal random effects for $T_{0}$ time points of channel $e$. Let $\bm{\Sigma}_{n}^{s}=\bm{V}_{n}\bm{R}_{n}^{s}\bm{V}_{n}^{\top}$, $\bm{R}_{n}^{s}$, and $\bm{R}_{n}^{t}$ be the spatial covariance matrix, spatial correlation matrix, and temporal correlation matrix for participant $n,n\in\{0,...,N\}$, respectively. $\bm{R}_{n}^{s}$ is a correlation matrix with the structure of compound symmetry characterized by a scale parameter $\eta_{n}\in(0,1)$; $\bm{V}_{k}$ is a diagonal matrix characterized by $(\sigma_{n,1},\cdots,\sigma_{\sigma_{n,E}})$, and $\bm{R}_{k}^{t}$ is a correlation matrix with the structure of exponential decay characterized by a scalar parameter $\rho_{n}\in(0,1)$. Given $Z_{n}=1$, where $n\in\{1,\cdots,N\}$ we assume that $\bm{\xi}_{n,i,j,t}$ and $\bm{\varepsilon}_{n,i,j,e}$ follow multivariate normal distributions with zero mean vectors and covariance matrices of the new participant; given $Z_{n}=0$, we assume that $\bm{\xi}_{n,i,j,t}$ and $\bm{\varepsilon}_{n,i,j,e}$ follow multivariate normal distributions with zero mean vectors and participant-specific covariance matrices. In summary,
@@
Let $\bm{\Theta}$ be a collection of all unknown parameters, where the primary parameters of interest are those associated with the new participant, i.e., $\bm{B}_{0,1},\bm{B}_{0,0},\sigma_{0,e}^{2},\eta_{0}$, and $\rho_{0}$. We are also interested in making inferences on the posterior probability of matching the new participant, i.e., $\operatorname{Pr}(Z_{n}=1\mid\bm{X},\bm{Y})$, for source participant $n$. The remaining parameters are considered as nuisance parameters.

### v2

@@ Abstract
An Event-Related Potential (ERP)-based Brain-Computer Interface (BCI) Speller System assists people with disabilities to communicate by decoding electroencephalogram (EEG) signals. A P300-ERP embedded in EEG signals arises in response to a rare, but relevant event (target) among a series of irrelevant events (non-target). Different machine learning methods have constructed binary classifiers to detect target events, known as calibration. The existing calibration strategy uses data from participants themselves with lengthy training time. Participants feel bored and distracted, which causes biased P300 estimation and decreased prediction accuracy. To resolve this issue, we propose a Bayesian signal matching (BSM) framework to calibrate EEG signals from a new participant using data from source participants. BSM specifies the joint distribution of stimulus-specific EEG signals among source participants via a Bayesian hierarchical mixture model. We apply the inference strategy. If source and new participants are similar, they share the same set of model parameters; otherwise, they keep their own sets of model parameters; we predict on the testing data using parameters of the baseline cluster directly. Our hierarchical framework can be generalized to other base classifiers with parametric forms. We demonstrate the advantages of BSM using simulations and focus on the real data analysis among participants with neuro-degenerative diseases.
Keywords: Bayesian Method, Mixture Model, Transfer Learning, P300, Brain-Computer Interface, Calibration-Less Framework.
@@ 1.1 Background
A Brain-Computer Interface (BCI) is a device that interprets brain activity to assist people with severe neuro-muscular diseases with normal communications. An electroencephalogram (EEG)-based BCI speller system is one of the most popular BCI applications, which enables a person to “type” words without using a physical keyboard by decoding the recorded EEG signals. The EEG brain activity has the advantages of non-invasiveness, low cost, and high temporal resolution (Niedermeyer and da Silva,, 2005). The P300 event-related potential BCI design, known as the P300 ERP-BCI design (Donchin et al.,, 2000), is one of the most widely used BCI frameworks and the focus of this work.
@@
An ERP is an event potential in the EEG signals in response to an external event. The P300 ERP is a particular ERP that occurs in response to a rare, but relevant event (i.e., highlighting a group of characters on the screen) among a group of frequent, but irrelevant ones. The relevant (target) P300 ERP usually has a positive deflection in voltage with a latency around 300 ms. Here, the latency refers to the time intervals between the onset of the event and the first response peak. Although P300 ERP is the most useful ERP, other types of ERPs, i.e., N200, exist and help interpret and classify brain activity. During the experiment, participants were asked to wear an EEG cap with multiple electrodes, sit next to a virtual screen, and focus on a specific character to spell, denoted as the “target key.” The study applied the row-and-column paradigm (RCP) developed by (Farwell and Donchin,, 1988), which is a 6 $\times$ 6 grid with 36 keys. Each event is either a row stimulus or a column stimulus. The order of rows and columns is random, but it loops through all rows and columns every 12 stimuli, known as a sequence. Participants were asked to mentally count whenever they saw a row or column stimulus containing the target key, and do nothing, otherwise. Thus, each sequence always contained exactly two stimuli that elicited target P300 ERPs.
@@
The left panel of Figure 1 demonstrates the fundamental mechanism of the P300 ERP-BCI design. The virtual screen presents a sequence of stimuli; Participants respond to the stimuli with EEG signals recorded; The machine truncates the EEG signals in a fixed time response window after each stimulus and generates a EEG feature vector. For raw data with multiple electrodes, the ultimate truncated signal segment is achieved by concatenating multi-channel EEG signals. The machine then makes a binary decision whether the participant’s brain elicits a target P300 ERP response. After stimulus-specific binary classifier scores are generated, they are converted into character-level probabilities. Finally, the character with the highest probability is selected and shown on the virtual screen. For offline spelling, the final decision per character is usually made after multiple sequence replications. Many machine learning (ML) methods have successfully constructed such binary classifiers for P300 ERPs, including stepwise linear discriminant analysis (swLDA) (Donchin et al.,, 2000), (Krusienski et al.,, 2008), support vector machine (SVM) (Kaper et al.,, 2004), independent component analysis (ICA) (Xu et al.,, 2004), linear discriminant analysis (LDA) with xDAWN filter (Rivet et al.,, 2009), convolutional neural network (CNN) (Cecotti and Graser,, 2010), and Riemannian geometry (RG) (Barachant et al.,, 2011). Recently, Ma et al., (2022) attempted to improve prediction accuracy by identifying spatial-temporal discrepancies between target and non-target P300 responses and explicitly addressed the overlapping ERP issue.
Figure 1: Left: The conventional offline framework for P300 ERP spelling includes signal acquisition, pre-processing and feature extraction, binary classification, character-level probability calculation, and final result demonstration. Right: The proposed data integration framework leverages data of source participants and small data from the new participant to construct the BSM binary classifier with participant-specific selection indicators.
1.2 Challenges and Existing Work
@@
The conventional framework for P300 ERP spelling task requires training and testing, known as calibration and free-typing, as well. Typically, participants were asked to copy a multi-character phrase to build a participant-specific training profile. Due to the low signal-to-noise ratio in EEG, current calibration strategy relies on collecting data from participants themselves with a fixed but large amount of sequence replications. However, participants tended to feel bored and tired after repetitions, leading to biased target P300 ERPs and inefficient calibration process. Therefore, it is important to identify efficient calibration methods while maintaining decent prediction accuracy for free-typing. Existing work has tackled this problem by applying transfer learning (Bozinovski and Fulgosi,, 1976). In statistical fields, we denoted this concept as data integration (Lenzerini,, 2002). To avoid confusion, we do not distinguish among “transfer learning,” “data integration,” or “borrowing” in this work. We briefly introduce existing works that have applied transfer learning under different domains with applications to P300 ERP-BCIs. A recent review (Wu et al.,, 2020) on transfer learning summarized methods between 2016 and 2020 with six paradigms and applications. For P300-ERP BCI paradigm, they considered cross-session/participant and cross-device transfer learning scenarios and produced a nice table for alignment-related approaches including offline/online status, known/unknown labels, alignment framework, referencing objects, base classifiers, and computational costs. We continue introducing the work that were not covered previously.
@@
Ensemble Learning Generic Information (ELGI) (Xu et al.,, 2015) examined whether inter-participant information is beneficial to ERP classification with simple base classifiers. They applied swLDA to each participant’s dataset as base classifier and combine them to construct the final classifier. Later, they introduced the Weighted Ensemble Learning Generic Information (WELGI) (Xu et al.,, 2016) by adding weights to each base classifier. Similarly, An et al.,in 2020 proposed a weighted participant-semi-independent classification method (WSSICM). They applied SVM and fit each base classifier by combining the entire data of each source participant and a small portion of data of the new participant. Weights for base classifiers were determined by an ad-hoc approach. Finally, Adair et al., proposed the Evolved Ensemble Learning Generic Information (eELGI) that developed a data-driven algorithm to determine the grouping criterion for training sets among source participants and applied swLDA to form base classifiers.
@@
Transfer learning based on Riemannian geometry has gained increasing attention recently due to its fast speed to converge and a natural framework to leverage information from source participants. The input data for RG were the sample covariance matrices, and the distance-based algorithm based on RG was called Minimum Distance to Mean (MDM) (Barachant et al.,, 2011). Congedo et al., and Barachant and Congedo, extended the original RG-based classifier to the P300 ERP design by augmenting the covariance matrix with cross-covariance between observed signals and reference target signals. Rodrigues et al., applied the Riemannian procrustes analysis (RPA) to address the heterogeneity of EEG signals across different sessions or participants. Before the authors applied the MDM classifier, they applied certain affine transformations to raw participant-level covariance matrices such that the resulting covariance matrices were less heterogeneous across sessions or participants while their Riemannian distances were preserved. Li et al., standardized the covariance matrices across participants by applying the affine transformation with the participant-specific Riemannian geometric mean covariance matrix. Finally, Khazem et al., proposed Minimum Distance to Weighted Mean (MDWM) by combining estimated mean covariance matrices from source participants and the new participant via the Riemannian distance with a trade-off hyper-parameter.
@@ 1.3 Our Contributions
To reduce the calibration time while maintaining similar prediction accuracy, we proposed a Bayesian Signal Matching (BSM) method to build a participant-dependent, calibration-less framework. BSM reduced the calibration time of a new participant by borrowing data from pre-existing source participants’ pool. We performed an intuitive clustering approach by simply borrowing data from eligible participants without data transformation and homogenization. BSM specified the joint distribution of stimulus-specific EEG signals from all participants via a Bayesian hierarchical mixture model. Our method has several advantages: (i). From the perspective of user experience, fewer training data required for calibration under this framework increased higher spelling efficiency and placed less burden on participants. (ii). From the perspective of inference, we performed pairwise selection between each source participant and the new participant and specified the baseline cluster corresponding to the new participant. The selection indicators represented how close source calibration data of each source participant resembled the data of the new participant. If the selection indicator was 1, we merged source participant’s data with the new participant’s data, otherwise, we kept the source participant’s data as a separate cluster. This approach did not focus on the similarity evaluation among source participants, and thereby, avoided the label-switching issue. (iii). From the perspective of prediction, we used the baseline cluster for directly prediction without refitting the model with the augmented data. The applied Bayesian framework for parameter estimation also “leveraged” the shared information from source participants implicitly and was robust to noises. Our proposed hierarchical framework is also flexible to generalize to other classifiers with parametric likelihood functions. (iv). From the perspective of computation, we performed the data analysis on the transformed ERP responses after xDAWN filter (Rivet et al.,, 2009), The xDAWN filter was a technique for dimensionality reduction. This pre-processing technique increased the signal-to-noise-ratio and matching probability, and improved the modeling efficiency compared to modeling the raw EEG signals or joint modeling transformed EEG signals.
The paper is organized as follows: Sections 2 and 3 present the framework for the Bayesian hierarchical mixture model and posterior inference, respectively. Sections 4 and 5 present the numerical results for simulation and real data analysis, respectively. Section 6 concludes the paper with a discussion.
@@ 2.1 Basic Notations and Problem Setup
2.2 The Framework of Bayesian Signal Matching
We made three assumptions to our proposed framework as follows: (i). We assume that P300 ERPs are the same regardless of target character to spell. To simplify, we let all source participants spell the same character $\omega$ with the same sequence replication size $I$, so we drop the target character index $l$ and the source participant index $n$ from $I_{n}$. The biological mechanism behind the speller system is the response to an unexpected stimulus. Under the RCP design, there were always two target stimuli supposed to elicit target P300 ERPs, and this rule applied regardless of the spelling characters. (ii). We assumed that target and non-target ERP functions shared the same covariance matrix within the same cluster. We simplified the temporal covariance matrix as a scaled AR(1) correlation matrix. Linear discriminant analysis (LDA) with its variations have been shown to be effective binary classifiers for this problem, we argued that the difference on the mean level had sufficient separation effect, and sharing the same covariance matrix reduced the number of parameters. Empirical correlation matrix also suggested that the correlation declined fast and they were similar for target and non-target stimuli, even after the xDAWN filter. The shared and simplified covariance matrix also reduced the number of parameters and made the proposed model parsimonious, especially when we considered borrowing data from source participants. (iii). We performed selection with respect to the target ERP data and we excluded non-target ERP data from the source participants’ pool. Despite a larger sample size of non-target ERP data, unfortunately, their signals showed fewer unique patterns and would add more noises to the model fitting. The right panel of Figure 1 provides a detailed illustration of the data integration framework of proposed Bayesian signal matching method. Details of three assumptions can be found in the supplementary material Section X.
@@ 2.2 The Framework of Bayesian Signal Matching
where $\bm{B}_{0,1},\bm{B}_{0,0}$, and $\bm{\epsilon}_{0,i,j}$ are target ERP matrix, non-target ERP matrix, and random noise matrix of the new participant, of the same size $T_{0}\times E$, respectively. For source participant $n,n>0$, we introduce a participant-specific binary indicator $Z_{n}\in\{0,1\}$ and assume that $\operatorname{Pr}(Z_{n}=1)=\pi_{n}$, where $\pi_{n}$ is the probability whether source participant $n$ matches the new participant. With $\{Z_{n}\}$, we consider a signal matching model for the EEG signal matrix for source participant $n$ as follows:
$\bm{X}_{n,i,j}$ $=\bm{B}_{n,1}Y_{n,i,j}+\bm{B}_{n,0}(1-Y_{n,i,j})+\bm{\epsilon}_{n,i,j},$ (2)
$\bm{B}_{n,1}$ $=\bm{B}_{0,1}Z_{n}+\tilde{\bm{B}}_{n,1}(1-Z_{n}).$
where $\bm{B}_{n,1},\tilde{\bm{B}}_{n,1},\bm{B}_{n,0}$, and $\bm{\epsilon}_{n,i,j}$ are target ERP matrix after matching, target ERP matrix before matching, non-target ERP matrix, and stimulus-specific random error matrix of source participant $n$, respectively. Equation  (2) assumes that each source participant has either has the new participant’s parameter set with a successful matching or their own parameter set. Note that we do not borrow non-target EEG data of source participants. The quantities $\bm{X}_{n,i,j},\tilde{\bm{B}}_{\cdot},\bm{B}_{\cdot}$, and $\bm{\epsilon}_{n,i,j}$ are matrices with the same size of $T_{0}$ by $E$.
@@
$\bm{\xi}_{n,i,j}$ $=(\bm{\xi}_{n,i,j,1},\cdots,\bm{\xi}_{n,i,j,T_{0}})^{\top},\quad\bm{\varepsilon}_{n,i,j}=(\bm{\varepsilon}_{n,i,j,1},\cdots,\bm{\varepsilon}_{n,i,j,E}),$
where $\bm{\xi}_{n,i,j,t},(t=1,\cdots,T_{0})$ is a vector of spatial random effects for $E$ channels at time point $t$, and $\bm{\varepsilon}_{n,i,j,e},(e=1,\cdots,E)$ is a vector of temporal random effects for $T_{0}$ time points of channel $e$. Let $\bm{\Sigma}_{n}^{s}=\bm{V}_{n}\bm{R}_{n}^{s}\bm{V}_{n}^{\top}$, $\bm{R}_{n}^{s}$, and $\bm{R}_{n}^{t}$ be the spatial covariance matrix, spatial correlation matrix, and temporal correlation matrix for participant $n,n\in\{0,...,N\}$, respectively. $\bm{R}_{n}^{s}$ is a correlation matrix with the structure of compound symmetry characterized by a scale parameter $\eta_{n}\in(0,1)$; $\bm{V}_{k}$ is a diagonal matrix characterized by $(\sigma_{n,1},\cdots,\sigma_{\sigma_{n,E}})$, and $\bm{R}_{k}^{t}$ is a correlation matrix with the structure of exponential decay characterized by a scalar parameter $\rho_{n}\in(0,1)$. Given $Z_{n}=1$, where $n\in\{1,\cdots,N\}$, we assume that $\bm{\xi}_{n,i,j,t}$ and $\bm{\varepsilon}_{n,i,j,e}$ are noise vectors generated by the new participant’s parameters; Given $Z_{n}=0$, we assume that $\bm{\xi}_{n,i,j,t}$ and $\bm{\varepsilon}_{n,i,j,e}$ are noise vectors generated by participant $n$’s parameters.
@@
Let $\bm{\Theta}$ be a collection of all unknown parameters. Our primary parameters of interest are those associated with the new participant, i.e., $\bm{B}_{0,1},\bm{B}_{0,0},\sigma_{0,e}^{2},\eta_{0}$, and $\rho_{0}$, and posterior probability of matching indicators, i.e., $\operatorname{Pr}(Z_{n}=1\mid\bm{X},\bm{Y}),n=1,\cdots,N$. The remaining parameters are considered as nuisance parameters.